2. Given a normal distribution with = 200 and o² = 100, find

(a) the area below 214;

(b) the area above 179;

(c) the area between 188 and 206;

(d) the x value that has 80% of the area below it;

(e) the two x values containing the middle 75% of the area.


1
Expert's answer
2021-07-05T15:14:59-0400

"\mu=200 \\

\sigma^2 = 100 \\

\sigma= 10"

(a)

"P(X<214) = P(Z< \frac{214-200}{10}) \\

=P(Z<1.4) \\

= 0.9192"

(b)

"P(X>179) = 1 -P(X<179) \\

= 1 -P(Z< \frac{179-200}{10}) \\

= 1 -P(Z< -2.1) \\

= 1 -0.01786 \\

= 0.98214"

(c)

"P(188<X<206) = P(X<206) -P(X<188) \\

=P(Z< \frac{206-200}{10}) -P(Z< \frac{188-200}{10}) \\

= P(Z< 0.6) -P(Z< -1.2) \\

= 0.72575 -0.11507 \\

= 0.61068"

(d)

"P(Z<0.8416) = 0.8 \\

\frac{x-200}{10}= 0.8416 \\

x = 200 + 10 \times 0.8416 \\

x= 208.416"

(e)

"P(-z<Z<z) = 0.75 \\

P(Z<z) -P(Z< -z) = 0.75 \\

P(Z<z) -(1 -P(Z<z)) = 0.75 \\

2P(Z<z) -1 = 0.75 \\

2P(Z<z) = 1.75 \\

P(Z<z) = 0.875"

The value of z that leaves an area of 0.875 to the right is z1 = -1.15 and the left is z2 = 1.15.

"-1.15 = \frac{x_1 -200}{10} \\

x_1 = 200 -1.15 \times 10 \\

x_1 = 188.5 \\

1.15 = \frac{x_2 -200}{10} \\

x_2 = 200 + 1.15 \times 10 \\

x_2 = 211.5"